# Characterizing and Efficiently Accelerating Multimodal Generation Model Inference

### v1

@@ III System Performance Characterization on Multi-Modal Model Inference
III-A Sequence Length and Latency Distribution
@@
Sequence length is a key task-specific dimension that determines where the most important performance acceleration opportunities come from. Sequence length distribution also affects the computational efficiency of generative models. For instance, models with shorter sequence lengths require less computation time to generate samples compared to models with longer sequence lengths. Transformer-based models, in particular, are highly sensitive to sequence length distributions due to their attention operation, where computational costs increase quadratically (i.e., $O(N^{2}d)$, where $N$ and $d$ denote sequence length and embedding dimension, respectively). In Table II, we delve into the sequence length distribution for the four different generative AI models.
@@
And in Figure 3, we show end-to-end inference latency distribution to show the correlation between the sequence length and the latency. We measure the inference latency of each sample with a batch size of 1 on NVIDIA A100 GPU to get the latency distribution. Based on our analysis, latency distribution is highly correlated to the sequence length distribution. We are going to discuss the correlation in detail in the following parts. By understanding the sequence length and latency distribution and its correlation, our goal is to better understand what determines the different system performance of four generative AI models and help optimizing those models by understanding the trade-off between the length of the sample and the computational efficiency.
@@
Llama: For the Llama-based coding tasks (Code Llama), we focus on programming problems and the evaluation of AI’s coding capabilities using the HumanEval [5] and MBPP [3], respectively. The input prompts describe the programming problems in text, such as ”Write a python function to find the first repeated character in a given string.”. We define input sequence lengths of Coda Llama as the number of text tokens fed into the model whereas output sequence lengths represent the number of text tokens generated by the model. In general, the input sequence length for MBPP is in the order of tens of tokens while the input sequence length for HumanEval is in the hundreds. This is because the HumanEval dataset gives more detailed constraints of problems with simple examples in the input prompts. In contrast, the output sequence lengths for HumanEval is in the order of hundreds since the solution for these datasets are quite simple that could be solvable in few lines of code (around 10 lines in general).
@@
In Table 3, we report the latency distribution of HumanEval and MBPP dataset. Overall, MBPP has longer end-to-end latency than HumanEval as the number of decoding step is the key factor in deciding the end-to-end latency which will be discussed in more detail in Section III-B Observation #1. Also, T-T tasks have the widest latency distribution among all tasks as the end-to-end latency has high correlation with the sequence lengths and the number of decoding steps distribution. The standard deviation is one of the most representative metric to show how broadly the values are distributed. Based on our analysis, T-T tasks have the largest standard deviation for input sequence lengths and decoding steps.
@@
Seamless: For sequence length analysis of Seamless, we focus on the Fleurs [7] dataset which contains the speech version of the FLoRes [17] machine translation benchmark in 102 different languages. This dataset is used for a variety of speech tasks, including automatic speech recognition (ASR), speech language identification, translation and retrieval.
@@
Chameleon: For Chameleon-based multi-modal tasks, we focus on the widely-used MSCOCO [23], Vizwiz [19] datasets for I-T&T-I and IT-T tasks, respectively.
@@
  • Image to Text (I-T) tasks: Chameleon uses newly trained image tokenizers [13] for the image input modality and the BPE tokenizer [36] for the text input modality. The image tokenizer generates 1024 image tokens per image. Thus, the I-T task has a fixed input sequence length of 1030 tokens, including the 1024 tokens representing an image and an additional 6 tokens representing the static prompting telling the model to generate the caption (e.g., ”Describe the figure”).
@@
  • Image/Text to Text (IT-T) tasks: For the IT-T generation task, a representative use case is Visual Question Answering (VQA), which generates response given an image and a question for the image, such as ”Can you tell me what this image is about?”. In this case, the input sequence is the concatenation of image tokens from the image tokenizer and text tokens from the text tokenizer, resulting in an input sequence length of 1024, plus the number of tokens representing the question. Taking the Vizwiz dataset as an example, the number of text tokens for the questions range from 3 to 65.
    Note that I-T and IT-T tasks have fixed number of decoding steps because Chameleon decodes until a maximum output length for a given task and then extracts the substring containing the desired predictions using task-specific templates. This is why the number of decoding steps remains fixed, even though the output lengths for different tasks may vary.
@@
    In Figure 3, I-T tasks generally have longer latency than IT-T tasks even though they have similar input sequence length. This is because the number of decoding steps of I-T tasks are longer than IT-T tasks. The maximum output length of I-T (30) is longer than that of IT-T (10) because generated text caption for image (image captioning) is generally longer than the generated text answers for the image (visual question answering, VQA) because VQA asks questions that can be answered in few words such as ”Q: Which one is the blue one? A: Right”, ”Q: ”What color is this A: White”.
@@
  • Text to Image (T-I) tasks: For the T-I generation task, instructions to generate image such as ”An upstairs living room is decorated nicely and holds a sewing machine.” is given as the text input prompt. Thus, the input sequence length is determined by the number of text tokens generated by the text tokenizer. We use the MSCOCO [23] dataset, for which the average input sequence length is 13.9.
@@
    In Figure 3, T-I tasks have the longest latency among the all tasks. Even though T-I tasks have shorter input sequence lengths, the number of decoding is highest which is 1024 resulting in the longest latencies. Also, as mentioned in Section II-A3, Chameleon uses a contrastive decoding methiod for T-I task, thus Chameleon runs the model twice at each incremental decoding step.
@@
gDLRM: For recommendation tasks with feature generation (HSTU), we focus on a synthetically generated dataset, where a sequence of user history is randomly generated. We generated 16384 number of inference samples, and the sequence of each sample comes with random integer indices, which range from 0 to 6000. The synthetically-generated sequence lengths are configured to represent the distribution we observed in the production environment as mentioned in the work [50].
@@
Even though the average input sequence length is 3000, the average of generated user history is 4813.9, because Also, as mentioned in Section II-A4, HSTU is composed of a stack of identical 14 layers but they limit the maximum input sequence length for later 11 layers as 1024 for speed improvement performance.
@@
Figure 4: Operator Time Breakdown of Code Llama [35], Seamless [6], Chameleon [42], HSTU [50]. P stands for the Prefill stage whereas D stands for the Decoding stage, respectively.
@@
III-B Operator Time Breakdown
@@
Figure 4 presents the end-to-end model inference time breakdown for Llama, Seamless, Chameleon and gDLRM. We characterize the inference time by maximizing the batch size for each workload to fit in the HBM memory capacity (i.e., 80GB) of a single NVIDIA A100 GPU [27]. We report the averaged breakdown result for 5 samples after 15 iterations of warmup for Code Llama, Seamless, Chameleon and HSTU (3 samples for T-I task of Chameleon model). For detailed description of the codebase and environment setup, please refer to Section IV.
Four generative AI models consist of different sets of operators. ”Idle” time indicates idle time on GPU device during the inference when GPU is in the idle status because of the GPU kernel launch overhead on CPU side. We divide prefill and decoding stage for Llama and Chameleon to better understand the different characteristics of each stage. And we show the normalized inference time to the end-to-end prefill time of Llama on top of each bar. Also note that we exclude embedding table lookup time of HSTU given that DLRM serving disaggregates embedding from the main model itself.
@@
For autoregressive generative models (Llama, Seamless, and Chameleon), the number of decoding steps matters the most to the end-to-end latency. As these models generate tokens sequentially, larger the number of decoding steps prolongs the generation process. For example, I-T task and IT-T task have similar average input sequence length while I-T task have 3 times higher number of decoding step according to Table II. This result in longer end-to-end inference latency as shown in Figure 4. Also, the T-I task in the Chameleon model takes the longest latency per inference sample because the image generation process involves 1024 decoding steps to produce a single image, significantly larger than the number of steps required by other tasks. This results in the longest latency per inference sample among the four models. Also, Llama has longer latency compared to I-T task and IT-T task of Chameleon even the input sequence lengths for Code Llama is much smaller (upto 13x). One of the primary cause for this is because Llama has higher number of decoding steps, resulting in the increased end-to-end latency.
@@
On the other hand, non-autoregressive models generate all tokens simultaneously rather than sequentially, so they can be significantly faster than autoregressive models during inference. This is particularly beneficial for long sequences or when real-time performance is crucial and this can lead to a better user experience. HSTU [50] demonstrates the potential benefits of non-autoregressive models.
@@
Seamless and HSTU have relatively higher GPU utilization compared to Llama and Chameleon. For Seamless, Speech/Text Encoder and Text Decoder are always activated and NAR T2U and Vocoder are selectively activated depending on the tasks. Among the four modules, only the text decoder is an autoregressive module indicating that only this module will be operating on matrix size of sequence length 1 except for Prefill while the rest modules are operating on the matrix size of the given full input sequence length. Thus, the overall GPU utilization for Seamless is higher than Llama and Chameleon since it has only one autoregressive module. For HSTU, the input sequence length is much larger (4813.9 x batch size) than other models according to Table II, so GPU spends much more time on computation resulting in high GPU utilization.
To address CPU-bound issue, optimizing techniques like torch.compile and CUDA Graph can significantly reduce the GPU kernel launch overhead. The latency improvement results and detailed explanations of torch.compile and CUDA Graph are provided in Section IV-A2.
@@
For Llama and Chameleon, Linear operation dominates the end-to-end inference time. For Seamless, linear operation takes a comparable portion of the inference time to attention operation. And for HSTU, attention operation dominates the inference time, unlike the other models. That is because the computation cost of attention operation grows quadratically ($O(N^{2})$) to the input sequence length and the input sequence length of HSTU is in higher order than other generative models as addressed in Table II.
Generally, the linear operation takes an insignificant amount of the inference time, so accelerating linear layer operations could bring more significant improvements to end-to-end latency than accelerating attention operations. In Section IV-B, we delve more deeply into inference acceleration using different numeric precision levels (e.g., FP16, BF16, or FP32) on the linear operation performance and output quality.
@@
Beam search maintains a beam of the K best sequences so far and considers the probabilities of the combination of all of the preceding words along with the word in the current position. Beam search maintains a separate copy of the KV cache for each sequence, and it needs to reorder KV caches for all attention layers according to the selected sequences from the previous decoding step to make sure each selected beam performs with the corresponding KV cache. This step requires copying all KV cache into a new memory space resulting in a significant portion of the inference runtime. This could be further optimized with torch.compile and we discuss the torch.compile case study for Seamless in Section IV-A2.
Model Code base Task Dataset Max. Batch Size # of Samples
Llama [4] T-T HumanEval 4 164
Chame- leon [42] I-T MSCOCO 16 5000
@@
T-I Coco Image 16 500
Seamless [24] S-S & S-T Fleurs 128 643
@@
HSTU [50] H-A Synthetic 32 16384
TABLE III: Datasets, codebase and batch size configuration for each workload.
Category Optimization Impact to Accuracy
System-level Optimization SDPA
torch.compile
AutoQuant
Workload-specific Optimization LayerSkip
CHAI
TABLE IV: Accuracy Impact of Optimization Techniques.
IV Accelerating Multi-Modal Model Inference via Cross-Stack Optimization
@@
  • System-level techniques optimize inference time performance horizontally across the generative AI tasks while being agnostic to specific algorithms. We consider Scaled Dot Product Attention (SDPA), torch.compile and CUDA graph optimization (CUDA Graph). SDPA leverages highly optimized and fused implementation to reduce the number of kernel launches and intermediate data transfers, which contributes to lower latency and memory usage. torch.compile and CUDA Graph facilitate streamline GPU task scheduling and execution, optimizing parallelism and resource utilization given system hardware. We also deploy quantization optimization using the PyTorch AutoQuant framework [30] — in Section IV-B. AutoQuant automates the parameter tuning process by determining the most efficient quantization method for each layer.
  • Workload-specific techniques optimize design objectives tailoring to algorithm or neural network specific characteristics. Taking a recent neural network optimization technique — LayerSkip [12] — tailor-designed for Transformer-based large language models, we customize and evaluate the impact of LayerSkip across the generative AI tasks in Section IV-C
Metholody Detail: Table III presents the datasets and the corresponding codebase used for each task, as well as the maximum batch size that fits in a single NVIDIA A100 GPU [27] used in our study. For the MSCOCO image dataset, we sub-sampled 500 out of 2000 data samples so the experiment time is more manageable and used the full dataset for the rest tasks. For HSTU, we generated synthetically a dataset with 16,384 samples, where a sequence of user history for each sample is randomly generated as explained in Section III-A. We validated and ensured the dataset is representative of production usecases.
IV-A Baseline is All You Need
IV-A1 Accelerating Attention

### v2

@@ 3 System Performance Characterization on Multi-Modal Model Inference
3.1 Sequence Length and Latency Distribution
@@
Sequence length is a key task-specific dimension that determines where the most important performance acceleration opportunities come from. Sequence length distribution also affects the computational efficiency of generative models. For instance, models with shorter sequence lengths require less computation time to generate samples compared to models with longer sequence lengths. Transformer-based models, in particular, are highly sensitive to sequence length distributions due to their attention operation, where computational costs increase quadratically (i.e., $O(N^{2}d)$, where $N$ and $d$ denote sequence length and embedding dimension, respectively). In Table 2, we delve into the sequence length distribution for the four different generative AI models.
@@
And in Figure 3, we show end-to-end inference latency distribution to show the correlation between the sequence length and the latency. We measure the inference latency of each sample with a batch size of 1 on NVIDIA A100 GPU to get the latency distribution. Based on our analysis, latency distribution is highly correlated to the sequence length distribution. We are going to discuss the correlation in detail in the following parts. By understanding the sequence length and latency distribution and its correlation, our goal is to better understand what determines the different system performance of four generative AI models and help to optimize those models by understanding the trade-off between the length of the sample and the computational efficiency.
@@
Llama: For the Llama-based coding tasks (Code Llama), we focus on coding capabilities of AI using the HumanEval Chen et al. (2021) and MBPP Austin et al. (2021), respectively. The input prompts describe the programming problems in text, such as "Write a python function to find the first repeated character in a given string.". We define input sequence lengths of Coda Llama as the number of text tokens fed into the model whereas output sequence lengths represent the number of text tokens generated by the model. In general, the input sequence length for MBPP is in the order of tens of tokens while the input sequence length for HumanEval is in the hundreds. This is because the HumanEval dataset gives more detailed constraints of problems with simple examples in the input prompts. In contrast, the output sequence lengths for HumanEval is in the order of hundreds since the solution for these datasets are quite simple that could be solvable in few lines of code (around 10 lines in general).
@@
In Figure 3, we report the latency distribution of HumanEval and MBPP dataset. Overall, MBPP has longer end-to-end latency than HumanEval as the number of decoding step is the key factor in deciding the end-to-end latency which will be discussed in more detail in Section 3.2 Observation #1. T-T tasks have the widest latency distribution among all tasks as the end-to-end latency has high correlation with the sequence lengths and the number of decoding steps distribution. The standard deviation is one of the most representative metric to show how broadly the values are distributed. Our observation says T-T tasks have the largest standard deviation for input sequence lengths and decoding steps.
@@
Seamless: For sequence length analysis of Seamless, we focus on the Fleurs Conneau et al. (2022) dataset which contains the speech version of the FLoRes Goyal et al. (2021) machine translation benchmark in 102 different languages. This dataset is used for a variety of speech tasks, including automatic speech recognition (ASR), speech language identification, translation and retrieval.
@@
Chameleon: For Chameleon-based multi-modal tasks, we focus on the widely-used MSCOCO Lin et al. (2015), Vizwiz Gurari et al. (2018) datasets for I-T&T-I and IT-T tasks, respectively.
@@
  • Image to Text (I-T) tasks: Chameleon uses newly trained image tokenizers Gafni et al. (2022) and the BPE tokenizer Sennrich et al. (2016) for the image and text input modality, respectively. The I-T task uses 1030 tokens, combining 1024 image tokens with 6 prompt tokens (e.g., "Describe the figure") for caption generation.
@@
  • Image/Text to Text (IT-T) tasks: For the IT-T generation task, a representative use case is Visual Question Answering (VQA), which generates response given an image and a question for the image, such as "Can you tell me what this image is about?". In this case, the input sequence combines 1024 image tokens with additional tokens for the question text. Taking the Vizwiz dataset as an example, the number of text tokens for the questions range from 3 to 65.
    Note that I-T and IT-T tasks maintain fixed decoding steps by using maximum output lengths and task-specific templates for prediction extraction.
@@
    In Figure 3, I-T tasks have longer output length (30) than IT-T (10) since image captioning require more words than VQA (visual question answering) answers, which are typically brief responses like "Q: Which one is the blue one? A: Right", "Q: "What color is this A: White".
@@
  • Text to Image (T-I) tasks: For the T-I generation task, instructions to generate image such as "An upstairs living room is decorated nicely and holds a sewing machine." is given as the text input prompt. Thus, the input sequence length is determined by the number of text tokens generated by the text tokenizer. We use the MSCOCO Lin et al. (2015) dataset, for which the average input sequence length is 13.9.
@@
    In Figure 3, T-I tasks have the longest latency among the all tasks. Even though T-I tasks have shorter input sequence lengths, the number of decoding is highest which is 1024 resulting in the longest latencies. Also, as mentioned in Section 2.1.2, Chameleon uses a contrastive decoding methiod for T-I task, thus Chameleon runs the model twice at each incremental decoding step.
@@
gDLRM: For recommendation tasks with feature generation (HSTU), we focus on a synthetically generated dataset, where a sequence of user history is randomly generated. We generated 16384 number of inference samples, and the sequence of each sample comes with random integer indices, which range from 0 to 6000. The synthetically-generated sequence lengths are configured to represent the distribution we observed in the production environment as mentioned in the work Zhai et al. (2025).
@@
Even though the average input sequence length is 3000, the average of generated user history is 4813.9, because Also, as mentioned in Section 2.1.4, HSTU is composed of a stack of identical 14 layers but they limit the maximum input sequence length for later 11 layers as 1024 for speed improvement performance.
@@
Figure 4: Operator Time Breakdown of Llama Rozière et al. (2024), Seamless Communication et al. (2023), Chameleon Team (2024a), HSTU Zhai et al. (2025). P stands for the Prefill stage whereas D stands for the Decoding stage, respectively.
@@
3.2 Operator Time Breakdown
@@
Figure 4 presents the end-to-end model inference time breakdown for Llama, Seamless, Chameleon and gDLRM. We characterize the inference time by maximizing the batch size for each workload to fit in the HBM memory capacity (i.e., 80GB) of a single NVIDIA A100 GPU NVIDIA (2020). We report the averaged breakdown result for 5 samples after 15 iterations of warmup for Code Llama, Seamless, Chameleon and HSTU (3 samples for T-I task of Chameleon model). For detailed description of the codebase and environment setup, please refer to Section 4.
Four generative AI models consist of different sets of operators. "Idle" time indicates idle time on GPU device during the inference when GPU is in the idle status because of the GPU kernel launch overhead on CPU side. We divide prefill and decoding stage for Llama and Chameleon to better understand the different characteristics of each stage. And we show the normalized inference time to the end-to-end prefill time of Llama on top of each bar. Note that we exclude embedding table lookup time of HSTU given that DLRM serving disaggregates embedding from the main model itself.
@@
For autoregressive generative models (Llama, Seamless, and Chameleon), the number of decoding steps matters the most to the end-to-end latency. As these models generate tokens sequentially, larger the number of decoding steps prolongs the generation process. For example, I-T task and IT-T task have similar average input sequence length while I-T task have 3 times higher number of decoding step according to Table 2. This result in longer end-to-end inference latency as shown in Figure 4. Also, the T-I task in the Chameleon model takes the longest latency per inference sample because the image generation process involves 1024 decoding steps to produce a single image, significantly larger than the number of steps required by other tasks. This results in the longest latency per inference sample among the four models. Also, Llama has longer latency compared to I-T task and IT-T task of Chameleon even the input sequence lengths for Llama is much smaller (upto 13x). One of the primary cause for this is because Llama has higher number of decoding steps, resulting in the increased end-to-end latency.
@@
On the other hand, non-autoregressive models generate all tokens simultaneously rather than sequentially, so they can be significantly faster than autoregressive models. This is particularly beneficial for long sequences or when real-time performance is crucial and this can lead to a better user experience. HSTU Zhai et al. (2025) demonstrates the potential benefits of non-autoregressive models.
@@
Seamless and HSTU have relatively higher GPU utilization compared to Llama and Chameleon. For Seamless, Speech/Text Encoder and Text Decoder are always activated and NAR T2U and Vocoder are selectively activated depending on the tasks. Among the four modules, only the text decoder is an autoregressive module indicating that only this module will be operating on matrix size of sequence length 1 except for Prefill while the rest modules are operating on the matrix size of the given full input sequence length. Thus, the overall GPU utilization for Seamless is higher than Llama and Chameleon since it has only one autoregressive module. For HSTU, the input sequence length is much larger (4813.9 x batch size) than other models according to Table 2, so GPU spends much more time on computation resulting in high GPU utilization.
To address CPU-bound issue, optimizing techniques like torch.compile and CUDA Graph can significantly reduce the GPU kernel launch overhead. The latency improvement results of torch.compile and CUDA Graph are provided in Section 4.1.2.
@@
For Llama and Chameleon, Linear operation dominates the end-to-end inference time. For Seamless, linear operation takes a comparable portion of the inference time to attention operation. And for HSTU, attention operation dominates the inference time, unlike the other models. That is because the computation cost of attention operation grows quadratically ($O(N^{2})$) to the input sequence length and the input sequence length of HSTU is in higher order than other generative models as addressed in Table 2.
Generally, the linear operation takes an insignificant amount of the inference time, thus accelerating linear layer operations could bring significant improvements to end-to-end latency than accelerating attention operations. In Section 4.2, we delve more deeply into inference acceleration using different numeric precision levels on the linear operation performance and output quality.
@@
Beam search maintains a beam of the K best sequences so far and considers the probabilities of the combination of all of the preceding words along with the word in the current position. Beam search maintains a separate copy of the KV cache for each sequence, and it needs to reorder KV caches for all attention layers according to the selected sequences from the previous decoding step to make sure each selected beam performs with the corresponding KV cache. This step requires copying all KV cache into a new memory space resulting in a significant portion of the inference runtime. This could be further optimized with torch.compile and we discuss the torch.compile case study for Seamless in Section 4.1.2.
Model& Codebase Task Dataset Max. Batch Size # of Samples
Llama Ben Allal et al. (2022) T-T HumanEval 4 164
Chame- leon Team (2024a) I-T MSCOCO 16 5000
@@
T-I Coco Img. 16 500
Seamless Meta (2023) S-S & S-T Fleurs 128 643
@@
HSTU Zhai et al. (2025) H-A Synthetic 32 16384
Table 3: Datasets, codebase and batch size configuration.
4 Accelerating Multi-Modal Model Inference via Cross-Stack Optimization
@@
  • System-level techniques optimize inference time performance horizontally across the generative AI tasks while being agnostic to specific algorithms. We consider Scaled Dot Product Attention (SDPA), torch.compile and CUDA graph optimization (CUDA Graph). SDPA leverages highly optimized and fused implementation to reduce the number of kernel launches and intermediate data transfers, which contributes to lower latency and memory usage. torch.compile and CUDA Graph facilitate streamline GPU task scheduling and execution, optimizing parallelism and resource utilization given system hardware. We deploy quantization optimization using the PyTorch AutoQuant framework PyTorch (2024) — in Section 4.2. It automates the tuning process by determining the most efficient quantization method for each layer.
  • Workload-specific techniques optimize design objectives tailoring to algorithm or neural network (NN) specific characteristics. Taking a recent NN optimization technique — LayerSkip Elhoushi et al. (2024) — tailor-designed for Transformer-based large language models, we evaluate the impact of LayerSkip across the generative AI tasks in Section 4.3
Metholody Detail: Table 3 presents the datasets and the corresponding codebase used for each task, as well as the maximum batch size that fits in a single NVIDIA A100 GPU NVIDIA (2020) used in our study. For the MSCOCO image dataset, we sub-sampled 500 out of 2000 data samples so the experiment time is more manageable and used the full dataset for the rest tasks. For HSTU, we generated synthetically a dataset with 16,384 samples, where a sequence of user history for each sample is randomly generated as explained in Section 3.1. We validated and ensured the dataset is representative of production usecases.
4.1 Baseline is All You Need
4.1.1 Accelerating Attention